# Right Reward Right Time for Federated Learning

### v2

@@ Abstract
Critical learning periods (CLPs) in federated learning (FL) refer to early stages during which low-quality contributions (e.g., sparse training data availability, strategic non-participation) can permanently impair the performance of the global model owned by the model owner (i.e., a cloud server). However, existing incentive mechanisms typically exhibit temporal homogeneity, treating all training rounds as equally important, thereby failing to prioritize and attract high-quality contributions during CLPs. This inefficiency is compounded by information asymmetry due to privacy regulations, where the cloud lacks knowledge of client training capabilities, leading to adverse selection. Thus, in this article, we propose a time-aware contract-theoretic incentive framework, named Right Reward Right Time (R3T), to encourage client involvement, especially during CLPs, to maximize the utility of the cloud server in FL. We formulate a cloud utility function that captures the trade-off between the achieved model performance and rewards allocated for clients’ contributions, explicitly accounting for client heterogeneity in time and system capabilities, effort, and joining time. Then, we devise a CLP-aware incentive mechanism deriving an optimal contract design that satisfies individual rationality, incentive compatibility, and budget feasibility constraints, motivating rational clients to participate early and contribute efforts. By providing the right reward at the right time, our approach can attract the highest-quality contributions during CLPs. Simulation and proof-of-concept studies show that R3T mitigates information asymmetry, increases cloud utility, and yields superior economic efficiency than conventional incentive mechanisms benchmarks. Our proof-of-concept results demonstrate up to a 47.6% reduction in the total number of clients and up to a 300% improvement in convergence time while achieving competitive test accuracy.
@@ I-A Background and Motivations
Federated learning (FL) is a privacy-preserving distributed learning paradigm, where distributed clients (e.g., devices or organizations) collaboratively train a shared artificial intelligence model under the coordination of a central server without disclosing private data [mcmahan2017communication, nguyen2025federated]. In classical FL settings, the importance of all training rounds and the efforts of distributed clients across these rounds are weighted equally. However, this setting is questioned by a recent finding in FL known as critical learning periods (CLPs), which highlight the varying impacts of different training rounds and significantly affect FL learning and training efficiency [yan2022seizing, huang2025pa3fed]. This finding was inspired by earlier research on CLPs in species’ cognitive and learning functions (i.e., humans and animals) and deep learning networks within a centralized setting [kandel2000principles, achille2018critical, jastrzebski2021catastrophic, golatkar2019time]. Particularly, CLPs are early learning phases (c.f., Fig. 1(1(a))) affecting long-term learning capabilities [hensch2005critical]. Empirical studies show that the absence of proper learning during these periods can lead to irreversible deficits. For instance, barn owls exposed to misaligned auditory and visual cues during their critical development periods fail to properly localize spatial locations in adulthood [knudsen1990sensitive]. This phenomenon was also observed in deep neural networks, in which any deficits and low effort (e.g., sparse data availability, late client participation) during this initial learning stage can cause lasting impairment on model performance, regardless of subsequent efforts (cf., Fig. 1(1(b))) [kleinman2023critical, yan2022seizing]. The reason for the phenomenon is still unclear, but possible explanations include the depth of the learning model, the structure of training data distribution, or defects in the implementation and training process [achille2018critical, jastrzebski2021catastrophic, golatkar2019time].
@@
Despite the promising merits of enhancing FL efficiency (e.g., accuracy and convergence rate), CLPs remain under-explored, including a lack of transparent, fair, and timely incentives for participation and contribution elicitation, and information asymmetry among clients and the global model owner (e.g., a cloud server [^1] The cloud server and the cloud are used interchangeably.) during these periods. Existing CLP-considered works in FL, such as CLPs detection, client selection, or CLP-aware aggregation, are based on a common assumption of voluntary engagement of clients in the training process (e.g., [yan2022seizing, yan2023criticalfl, huang2025pa3fed, 11060892]). However, this is impractical because clients are rational, strategic, benefit-driven, and need sufficient rewards to compensate for their efforts (e.g., data, computation resources, and time costs). This leads to a lack of client motivation for contribution, resulting in the unsustainable development of FL systems. Although several works have been proposed to bring incentivization by considering effort-reward approaches to fairly compensate FL clients’ effort (e.g., [kang2019incentive, 10097767, chen2024fdfl, 10323190]), they treat clients’ efforts and their impacts on the global model performance on every training round equally and ignore the temporal factor, whereas efforts in CLPs have more weights and mainly contribute to the overall performance of the global model. By failing to prioritize CLP-driven incentive solutions, current methods inefficiently allocate cloud rewards (e.g., money), thereby reducing cloud utility—defined as the difference between gains from global model performance and compensation costs (which is defined later in Section II).
@@
Furthermore, the design of efficient incentives is hindered by information asymmetry between the FL clients and the cloud server. Specifically, due to privacy restrictions, the cloud lacks prior knowledge of the clients’ private training capabilities, including the data, communication, and computation resources, efforts (e.g., data quantity and distribution), or available joining time. Therefore, assumptions that the cloud server knows precise information [10539084, zhan2020learning, sarikaya2019motivating], such as clients’ computing resources or clients’ data size, may not hold in practice. A compounded challenge overlooked by prior works is that existing contract-theoretic incentive approaches, while effective for handling information asymmetry in FL incentives (e.g., [sun2021pain]), fail to incorporate the temporal heterogeneity of CLPs, where early deficits cause irreversible model impairments regardless of later efforts [yan2022seizing, huang2025pa3fed]. This leads to suboptimal contract designs that treat all training rounds equally, inefficiently allocating incentives without prioritizing high-quality contributions [^2] We refer to high-quality contributions to high training efforts, such as a huge amount of closer-distribution-to-the-learning-task data and large computation resources, and they are available in CLPs. during CLPs, resulting in diluted contributions and permanent performance losses under asymmetry about clients’ training capabilities.
@@
There is also uncertainty regarding whether clients will exert the promised effort post-contract [bolton2004contract]. Even with contractually specified incentives, clients may engage in moral-hazard behaviors (e.g., low-effort training or strategic dropouts), and the cloud may be accused of delaying or reneging on payments, which can lead to disputes. Addressing these issues requires costly monitoring and verification and remains error-prone, increasing the difficulty of sustaining FL and implementing fair incentive mechanisms [10097767]. In this article, we aim to address the following research question (RQ).
@@
RQ: How can we design an incentive mechanism to attract rational clients with highest-quality contributions to improve FL learning and training efficiency, and mitigate information asymmetry between clients and the cloud, given the crucial role of CLPs in determining the final FL model performance?
@@
To address the aforementioned challenges, we propose a CLP-aware incentive contract-theoretic framework, named Right Reward Right Time (R3T). Leveraging contract theory, R3T effectively mitigates the information asymmetry issues by designing a set of temporally integrated contract items for the cloud server [bolton2004contract]. The optimal contract design incentivizes clients to self-reveal their training capabilities, join early, and contribute efforts by optimally prioritizing and providing corresponding generous rewards in the early training phase under complete and incomplete information cases, while satisfying individual rationality, incentive compatibility, and budget feasibility constraints. By providing the right reward at the right time, R3T enables the highest-quality clients to join early and contribute efforts during CLPs. R3T also integrates with blockchain smart contracts to handle the issues of centralized incentive governance and disputes caused by relying on the cloud server [wang2022infedge, yu2023ironforge, 3718082].
I-B Related Work
In the following, we discuss related work concerning CLPs in artificial neural networks and incentive mechanisms.
Critical learning periods in neural networks. Achille et al. in [achille2018critical] empirically investigated the correlation between critical periods in the memorization phase and the training outcome in deep neural networks. Their findings demonstrated that, analogous to animals, sensory deficits in CLPs cause lasting impairments to the information and connectivity of the network, as quantified by an approximation of the Fisher Information, and adversely affect learning outcomes regardless of subsequent additional training. Similarly, Kleinman et al. in [kleinman2024critical] analytically found that the depth of the deep neural network model and data distribution structure are fundamental sources of CLPs in deep networks, not only being explained by ascribing to biological aging processes observed in animals. A similar pattern was found in the distributed training manner, where Yan et al. in [yan2022seizing, yan2023criticalfl] showed that FL exhibits CLPs, which play a vital role in determining the final training performance, given the complexity of statistical and system heterogeneity. To further explain the phenomenon, metrics such as federated Fisher Information matrix and CLP-aware federated norm vector are proposed to trace and detect CLPs.
However, these CLP-aware FL studies assume voluntary participation and overlook costs and benefit-driven and strategic behavior, which do not reflect real-world conditions. Without proper CLP-target incentives, it may lead to the lack of high-quality contributions in CLPs that have been shown to be permanent and irrecoverable impairment on the global model performance, despite subsequent efforts.
Incentive mechanisms. By incorporating incentive mechanisms into FL systems, engagement and cooperation are promoted where clients are rational, resulting in enhanced efficiency [anonymous2026incentives]. Realizing the importance of incentive mechanisms, there are several efforts to develop game-theoretic incentive techniques in FL (e.g., [10264195, 9497718, han2024dynamic, sun2021pain, ding2023joint, 10557146, 10418498]). Authors in [10264195, 9497718, han2024dynamic] designed leader-follower-based incentive mechanisms to encourage clients’ contribution, where the cloud server sets the total reward or price per contribution unit, and then FL clients determine contribution strategies (e.g., data quantity or local performance) to maximize their utilities. However, the information asymmetry condition was not considered in these studies. To handle this issue, utilizing the self-revelation principle of contract theory [bolton2004contract], Sun et al. in [sun2021pain] proposed an incentive mechanism using a single client’s property (i.e., privacy leakage costs) to create different contract items under both complete and incomplete information cases to compensate for client’s privacy leakage while ensuring comparable learning model performance. Under the same information cases, Huang et al. in [10557146] proposed a data quality-based incentive mechanism, where the client’s data size, data quality, and unit data computing cost are considered, to derive optimal contract items to optimize the cloud server’s costs in the era of artificial intelligence-generated. Tang et al. in [10418498] proposed a non-cooperative game-based incentive mechanism to model the interactions between clients (i.e., organizations), deriving optimal processing capacity of each client, to address free-rider problems of public goods in FL under constraints of individual rationality and budget balance.
To provide further comprehensive incentive solutions by handling the transparency, centralization, and potential disputes among clients and the cloud server, blockchain is fused with FL systems to manage incentive calculation and distribution [wang2023incentive, han2024dynamic, yuan2023tradefl, 10418498]. For instance, Yuan et al. in [yuan2023tradefl] utilized Ethereum-based smart contracts to autonomously execute an incentive mechanism, that considers influences from competition and provides fair and non-repudiative compensation and incentive, to improve social welfare.
Nevertheless, existing works on incentive mechanisms in FL to provide the right reward at the right time have not been explored yet. None of these works consider the existence of CLPs, which have been shown to play a vital role in determining final FL model performance, and these studies treat every training round as equally important. Besides, information asymmetries among clients and the cloud server, and their impacts on CLPs’ advantages and the final model performance have not been analyzed yet.
To the best of our knowledge, R3T is the first CLP-aware incentive mechanism, considering the joint importance of joining time and effort, to attract the highest-quality rational clients to participate in the FL model training in early stages and mitigate information asymmetry between clients and the cloud server. R3T provides a transparent and autonomous execution of reward calculation and distribution.
I-C Contributions
@@
II System model
II-A System Overview

### v3

@@ Abstract
Critical learning periods (CLPs) in federated learning (FL) refer to early stages during which low-quality contributions (e.g., sparse training data availability, strategic non-participation) can permanently impair the performance of the global model owned by the model owner (i.e., a cloud server). However, existing incentive mechanisms typically assume temporal homogeneity, treating all training rounds as equally important, thereby failing to prioritize and attract high-quality contributions during CLPs. This inefficiency is compounded by information asymmetry due to privacy regulations, where the cloud lacks knowledge of client training capabilities, leading to adverse selection and moral hazard. Thus, in this article, we propose a time-aware contract-theoretic incentive framework, named Right Reward Right Time (R3T), to encourage client involvement, especially during CLPs, to maximize the utility of the cloud server in FL. We formulate a cloud utility function that captures the trade-off between the achieved model performance and rewards allocated for clients’ contributions, explicitly accounting for client heterogeneity in time and system capabilities, effort, and joining time. Then, we devise a CLP-aware incentive mechanism deriving an optimal contract design that satisfies individual rationality, incentive compatibility, and budget feasibility constraints, motivating rational clients to participate early and contribute efforts. By providing the right reward at the right time, our approach can attract the highest-quality contributions during CLPs. Simulation and proof-of-concept studies show that R3T mitigates information asymmetry, increases cloud utility, and yields superior economic efficiency compared to conventional incentive mechanisms. Our proof-of-concept results demonstrate up to a 47.6% reduction in the total number of clients and up to a 300% improvement in convergence time while achieving competitive test accuracy.
@@ I-A Background and Motivations
Federated learning (FL) is a privacy-preserving distributed machine learning (ML) paradigm, where distributed clients (e.g., devices or organizations) collaboratively train a shared artificial intelligence model under the coordination of a central server without disclosing private data [mcmahan2017communication, nguyen2025federated]. In classical FL settings, the importance of all training rounds and the efforts of distributed clients across these rounds are weighted equally. However, this setting is questioned by a recent finding in FL known as critical learning periods (CLPs), which highlight the varying impacts of different training rounds and significantly affect FL learning and training efficiency [yan2022seizing, huang2025pa3fed]. This finding was inspired by earlier research on CLPs in species’ cognitive and learning functions (i.e., humans and animals) and deep learning networks within a centralized setting [kandel2000principles, achille2018critical, jastrzebski2021catastrophic, golatkar2019time]. Particularly, CLPs are early learning phases (c.f., Fig. 1(1(a))) affecting long-term learning capabilities [hensch2005critical]. Empirical studies show that the absence of proper learning during these periods can lead to irreversible deficits. For instance, barn owls exposed to misaligned auditory and visual cues during their critical development periods fail to properly localize spatial locations in adulthood [knudsen1990sensitive]. This phenomenon was also observed in deep neural networks, in which any deficits and low effort (e.g., sparse data availability, late client participation) during this initial learning stage can cause lasting impairment on model performance, regardless of subsequent efforts (cf., Fig. 1(1(b))) [kleinman2023critical, yan2022seizing]. The reason for the phenomenon is still unclear, but possible explanations include the depth of the learning model, the structure of training data distribution, or defects in the implementation and training process [achille2018critical, jastrzebski2021catastrophic, golatkar2019time].
@@
Despite the promising merits of enhancing FL efficiency (e.g., accuracy and convergence rate), CLPs remain under-explored, including a lack of transparent, fair, and timely incentives for participation and contribution elicitation, and information asymmetry among clients and the global model owner (e.g., a cloud server [^1] The cloud server and the cloud are used interchangeably.) during these periods. Existing CLP-considered works in FL, such as CLPs detection, client selection, or CLP-aware aggregation, are based on a common assumption of voluntary engagement of clients in the training process (e.g., [yan2022seizing, yan2023criticalfl, fedprime, huang2025pa3fed, 11060892]). However, this is unrealistic in practice because clients, acting as rational, strategic, and benefit-driven agents, participate only when they receive adequate compensation for their efforts (e.g., data, computation resources, time costs, and potential privacy leakage risks). Moreover, the learned global model is considered a non-excludable public good, meaning that clients benefit from the model regardless of their individual effort levels [10418498, anonymous2026incentives]. Consequently, this may incentivize free-riding and weaken truthful efforts, which can degrade global model quality. The resulting performance deterioration can, in turn, impair service quality and user experience for downstream consumers of the model (e.g., data and model marketplaces [9025575]), threatening the long-term sustainability of FL deployments and services. Although several works have been proposed to bring incentivization by considering effort-reward approaches to fairly compensate FL clients’ effort (e.g., [anonymous2026incentives, murhekar2024you, kang2019incentive, 10097767, chen2024fdfl, 10323190]), they treat clients’ efforts and their impacts on the global model performance on every training round equally and ignore the temporal factor, whereas efforts in CLPs have more weights and mainly contribute to the overall performance of the global model. By failing to prioritize CLP-driven incentive solutions, current methods inefficiently allocate cloud rewards (e.g., money), thereby reducing cloud utility—defined as the difference between gains from global model performance and compensation costs (cf. Section III).
@@
Furthermore, the design of efficient incentives is hindered by information asymmetry between the FL clients and the cloud server. Specifically, due to privacy restrictions, the cloud lacks prior knowledge of the clients’ private training capabilities, including the data, communication, and computation resources, efforts (e.g., data quantity and distribution), or available joining time. Therefore, assumptions that the cloud server knows precise information [10539084, zhan2020learning, sarikaya2019motivating, murhekar2023incentives], such as clients’ computing resources or clients’ data size, may not hold in practice. A compounded challenge overlooked by prior works is that existing contract-theoretic incentive approaches, while effective for handling information asymmetry in FL incentives (e.g., [sun2021pain]), fail to incorporate the temporal heterogeneity of CLPs, where early deficits cause irreversible model impairments regardless of later efforts [yan2022seizing, huang2025pa3fed]. This leads to suboptimal contract designs that treat all training rounds equally, inefficiently allocating incentives without prioritizing high-quality contributions [^2] We refer to high-quality contributions to high training efforts, such as a huge amount of closer-distribution-to-the-learning-task data and large computation resources, and they are available in CLPs. during CLPs, resulting in diluted contributions and permanent performance losses under asymmetry about clients’ training capabilities.
@@
There is also uncertainty regarding whether clients will exert the promised effort post-contract [bolton2004contract]. Even with contractually specified incentives, clients may engage in moral-hazard behaviors (e.g., low-effort training or strategic dropouts), and the cloud may be accused of delaying or reneging on payments, which can lead to disputes. Addressing these issues requires costly monitoring and verification and remains error-prone, increasing the difficulty of sustaining FL and implementing fair incentive mechanisms [10097767]. Motivated by the aforementioned research gaps and challenges, in this article, we aim to address the following research question (RQ).
@@
RQ: How can we design an incentive mechanism to attract rational and benefit-driven clients with the highest-quality contributions to improve FL learning and training efficiency, while simultaneously mitigating information asymmetry between clients and the cloud, given the crucial role of CLPs in determining the final FL model performance?
@@
To answer the RQ, we propose a CLP-aware contract-theoretic incentive framework, named Right Reward Right Time (R3T). Leveraging contract theory, R3T effectively mitigates the information asymmetry issues by designing a set of temporally coupled contract items offered by the cloud server [bolton2004contract]. The resulting optimal contract induces clients to self-reveal their training capabilities, join early, and exert efforts by assigning higher priority and appropriately larger rewards to early-stage participation under both complete and incomplete information cases, while satisfying individual rationality, incentive compatibility, and budget feasibility constraints. By delivering the right reward at the right time, R3T attracts the highest-quality clients to join early and contribute efforts during CLPs, where these efforts are most consequential for final model performance. R3T also integrates with blockchain smart contracts to decentralize incentive governance and reduce disputes caused by relying solely on the cloud [wang2022infedge, yu2023ironforge, 3718082].
I-B Contributions
@@
The remainder of this paper is organized as follows. Section II reviews the related work, followed by the system model in Section III. Sections IV and V present the proposed game-theoretic R3T framework and its experimental results, respectively. Section VI concludes this paper.
II Related Work
This section reviews relevant background and advances in CLPs for artificial neural networks and incentive mechanisms for FL. The intersectional gap between these domains motivates our research contribution. The comparison of prior research and this work is shown in Table I.
II-A Critical Learning Periods in Centralized ML and FL
A significant amount of research has examined CLPs, including their effects on training efficiency and learning generalization, their interactions with optimization hyperparameters, and methods to detect and exploit them in both centralized and distributed learning settings [achille2018critical, kleinman2024critical, fukase2025one, yan2022seizing, yan2023criticalfl, fedprime, yan2023defl, 11302878, 11275875, 10802950, ecolearnclps, 10737242, li2023revisiting].
In centralized deep learning, Achille et al. in [achille2018critical] empirically linked CLPs, especially in the memorization phase, to the downstream performance. Their findings demonstrated that, analogous to animals, sensory deficits in CLPs cause lasting impairments to the learning model network’s information and connectivity, as quantified by an approximation of Fisher Information, and adversely affect learning outcomes, regardless of subsequent additional training. Complementing this empirical evidence, Kleinman et al. in [kleinman2024critical] provided analytical insights indicating that the deep neural network model depth and the structure of the data distribution are fundamental sources of CLPs in deep networks, not solely explained by biological aging processes observed in animals. More recently, Fukase et al. in [fukase2025one] studied a practical problem of identifying CLPs and proposed a detection approach based on the cosine distance between the initial and final model weights.
TABLE I: Comparison between prior research & R3T.
[…]
Work Venue Year F1 F2 F3 F4 F5
Critical Learning Periods in Centralized and FL
[achille2018critical] ICLR 2018 $Circle \CIRCLE \Circle \Circle$
[yan2022seizing] AAAI 2022 $Circle \CIRCLE \Circle \Circle$
[li2023revisiting] ICML 2023 $Circle \CIRCLE \Circle \Circle$
[yan2023criticalfl] SIGKDD 2023 $Circle \CIRCLE \Circle \Circle$
[fedprime] PKDD 2024 $Circle \CIRCLE \Circle \Circle$
[11302878] TMC 2025 $Circle \CIRCLE \CIRCLE \Circle$
[11275875] TMC 2025 $Circle \CIRCLE \Circle \Circle$
[10737242] TMC 2025 $Circle \CIRCLE \Circle \Circle$
Incentive Mechanisms for FL
[…]
[10539084] TNSE 2024 $CIRCLE \Circle \Circle \LEFTcircle$
[…]
[11363323] TNSM 2026 $CIRCLE \Circle \Circle \Circle$
[anonymous2026incentives] ICLR 2026 $CIRCLE \Circle \Circle \Circle$
R3T - 2026 $CIRCLE \CIRCLE \CIRCLE \CIRCLE$
F1: Incentivization; F2: Information asymmetry; F3: CLP awareness; F4: Proof of concept; F5: decentralized incentive management.
Parallel findings have been reported in FL. Yan et al. in [yan2022seizing, yan2023criticalfl] showed that FL also exhibits CLPs that disproportionately shape final model performance under data statistical and system heterogeneity. To characterize and detect these phases, they introduced CLP-tracking metrics such as the federated Fisher Information matrix and CLP-aware federated norm vector. Building on these foundations, subsequent studies [fedprime, ecolearnclps, 11275875] leveraged CLP awareness for system-level optimization. For instance, CLP-guided client selection has been developed using local client utilities that capture factors such as energy consumption, estimated contribution effort to the global model, and a tailored CLPdetection has been proposed to selectively upload updates when selected clients are inferred to operate within CLPs [fedprime]. Furthermore, authors in [10737242, 11302878] extended CLP-aware designs to asynchronous and mobile FL by dynamically scheduling updates and adjusting training hyperparameters to mitigate global model generalization degradation caused by test-data shifts and temporal class imbalances. Li et al. in [li2023revisiting] improved the global model generalization and mitigated heterogeneity by studying the link of CLPs and their proposed adaptive aggregation, which upweights more coherent clients during early training and stops the learned reweighting around the end of CLPs.
Despite these advancements, the majority of CLP-aware FL studies implicitly assume voluntary participation and largely overlook costs and benefit-driven and strategic behavior of rational clients. In practical FL deployments, clients may withhold data, communication, and computational resources, or strategically degrade model update quality, if participation costs are not offset by adequate incentives. Consequently, without explicit CLP-targeted incentive mechanisms, the FL systems may fail to elicit high-quality contributions required during CLPs, which prior evidence has shown to result in long-term irrecoverable impairment on the global model performance, despite subsequent efforts.
II-B Incentive Mechanisms for Federated Learning
Incentive mechanisms have been widely incorporated into FL to stimulate participation and sustain cooperation among rational and benefit-driven clients, thereby improving convergence efficiency and the resulting global model quality [10323190, anonymous2026incentives, 10739914, 10539084, 10264195, 9497718, han2024dynamic, sun2021pain, ding2023joint, 10557146, 10418498, wu2024incentive, 11363323, murhekar2023incentives, 103714961, 10586269, chen2024fdfl, yuan2023tradefl, le2025applications].
A dominant line of work adopts game-theoretic formulations, which provide a principled framework for modeling strategic interactions and utility-maximizing decisions of distributed clients and the cloud server in FL (e.g., [10264195, 9497718, han2024dynamic, sun2021pain, ding2023joint, 10557146, 10418498]). Specifically, authors in [ding2023joint, 103714961, 10323190, 10264195, 9497718, han2024dynamic, 10586269, 10739914] designed leader-follower-based incentive mechanisms to encourage contribution in one and multiple training rounds, where the cloud server first announces the total reward or price per contribution unit, and then clients determine effort strategies (e.g., data volume or computing power for training global model or cleaning noisy local data samples) to maximize individual utilities and system-wide social welfare [murhekar2023incentives]. However, many of these formulations rely on simplified information assumptions and do not explicitly model server-client information asymmetries, which is exacerbated by data-privacy regulations that restrict the disclosure of clients’ local data characteristics (e.g., data volume, quality, and privacy costs). To handle incomplete information, contract-theoretic designs exploit the self-revelation principle [bolton2004contract]. For instance, Sun et al. in [sun2021pain] constructed contract items based on privacy-leakage costs under both complete and incomplete information to compensate privacy loss while preserving learning performance, and Xu et al. in [11363323] further improved contract-based schemes by differentiating client contributions in the case clients have the same contract type, thereby enhancing the fairness. Under the same information conditions, Huang et al. in [chen2024fdfl, 10557146, anonymous2026incentives, 10539084, 10586269] proposed contribution-based incentive mechanisms, where the client’s data size and/or data quality are incorporated, to derive optimal clients’ contributions that minimize the cloud server’s costs. Tang et al. in [10418498] proposed a non-cooperative game-based incentive mechanism to model the interactions between clients (i.e., organizations), deriving optimal processing capacity of each client, to address free-rider problems of public goods in FL under constraints of individual rationality and budget balance. Wu et al. in [wu2024incentive] targeted problems of reward timeliness and monetary reward infeasibility in FL incentivization by providing a proportion of local model updates as rewards proportional to the client’s per-round contributions.
In order to improve transparency and reduce centralization-induced disputes, blockchain has also been integrated with FL to support auditable incentive computation and automated distribution [wang2023incentive, han2024dynamic, yuan2023tradefl, 10418498, 10539084, nguyen2021federated]. For instance, Yuan et al. in [yuan2023tradefl] leveraged Ethereum smart contracts to autonomously execute an incentive mechanism that explicitly accounts for competitive effects and provides fair and non-repudiative compensation, to enhance social welfare.
Nevertheless, existing incentive designs largely treat all training rounds as equally valuable and, consequently, do not yet address how to deliver the right reward at the right time. In particular, none of these works considers the presence of CLPs, which have been shown to shape final FL model performance; nor do they analyze how information asymmetries between clients and the cloud server interact with CLPs to influence incentive effectiveness and final learning outcomes.
To the best of our knowledge, R3T is the first CLP-aware incentive mechanism that jointly accounts for client joining time and contribution effort, to recruit rational and benefit-driven clients with the highest-quality contributions during early learning stages, while mitigating information asymmetry between clients and the cloud server. R3T enables transparent and autonomous reward computation and distribution.
III System model
III-A System Overview